3. A 950-kg cylindrical can buoy floats vertically in salt water. The diameter of the buoy is 0.900m. Calculate the additional distance the buoy will sink when a 70.0-kg man stands on top of it.

1
Expert's answer
2022-06-14T08:42:18-0400

"\rho gV=mg,"

"\rho\frac{\pi d^2}4h=m,"

"h=\frac{4m}{\pi\rho d^2}=10~cm."